### v2

@@ 5.0.1 Availability and Quality
Building and employing deep learning models face several challenges. The training of deep learning requires a large number of instances (examples) to achieve high accuracy and generalization Munappy et al. (2022). Furthermore, the complexity of deep neural networks may lead to overfitting, where the model performs well on training data but fails to generalize on new, unseen data. This phenomenon frequently arises when the models is trained on insufficient data, highlighting the importance of diverse and extensive datasets. However, the data collection and annotation are time consuming and often require domain experts, specialized training and standardization Luca et al. (2022). Moreover, this process is prone to error and has the risk of introducing biases into the dataset which can significantly impact the performance of the trained model. One of the approaches to address this issue is transfer learning. Transfer learning involves the use of a deep learning model (known as pre-trained model) that is trained on a large dataset for solving a specific task (with a small dataset) Zhuang et al. (2020). The pre-trained model serves as a basis for the model training by fine-tuning the weights of the pre-trained model and adapting it to the new prediction task. This approach helps to mitigate the lack of training data in the target domain. Furthermore, transfer learning reduces computational resources required to train the model and helps faster convergence.
@@
Another approach that can be employed to address the lack of data is data augmentation. Data augmentation is a convenient method that increases the number of instances by performing transformation functions on the existing instances without changing the labels Mumuni and Mumuni (2022). In the domain of computer vision, image transformation such as rotation, translation and cropping. However, it is important to consider the output of the transformation because the resultant may not represent the actual data. For example, flipping or adding noise to a signal may introduce distortion or changing the characteristics (trend, seasonality and cyclic variations) of the signal. Thus, careful consideration must be given to ensure that the generated instances still accurately represent the underlying patterns present in the data. Data augmentation can also be realized by generating synthetic data to supplement the training set. Synthetic data is artificially created data that resembles real data but is generated using statistical methods or deep generative models Hu et al. (2023); Murtaza et al. (2023). The generated data can complement the less-diverse, limited datasets, providing a broader range of examples for the model to learn from. However, generating synthetic data that accurately reflects the characteristics of the real-world data is challenging. Careful consideration must be given to the choice of models and parameters used to ensure the synthetic data is realistic and representative of the real-world data.
5.0.2 Interpretability and Explainability
@@
Interpretability and explainability is crucial for building trust and understanding how predictive models make decisions especially in high-stake applications such as healthcare and medical image analysis Tonekaboni et al. (2019). However, as deep learning models become more intricate and complex with numerous layers, subnetworks and a large number of parameters, the models are often perceived as a “black box” and difficult to explain in terms of decision-making processes. Therefore, it is crucial for the researchers to focus on methods that provide insights into how a deep learning model performs the prediction and how its decisions are influenced by the input data, making it more transparent and trustworthy. Numerous methods have been proposed for interpreting and explaining the decisions of deep learning models which can be categorized into visualization (feature attribution), model distillation and intrinsic (explainable by itself). Visualization methods involve the use of scientific visualization such as saliency maps or heatmaps to express the explanation by highlighting the degree of association between the inputs and the predictions Tjoa et al. (2023). The heatmaps identify the saliency of the input features influencing the model’s predictions. The visualization approach is simple and intuitive and can be applied to tabular data and image data. Furthermore, it can be used to identify and debug issues in deep learning models, leading to improved performance and robustness.
@@
Model distillation is an approach to approximating a complex model by fitting a simpler model using the training set. The simpler model is built typically using a simpler or interpretable algorithm such as linear regression, decision tree or rule-based methods Li and Shen (2024). In this approach, the simpler model is trained to resemble the predictive behavior of the complex model. Then, the simpler model may serve as the proxy or surrogate model for explaining the complex model. Model distillation can be used together with visualization to further enhance the interpretability of the complex model Termritthikun et al. (2023). Model distillation seeks explanations of the models that were never designed to be explainable. Ideally, the explanation of a deep learning model’s prediction should be included as part of the model output, or the explanation can be derived from the architecture of the model. This is because an intrinsic model can learn not only the mapping between the input and output, but also generate an explanation of the prediction that is faithful to the model’s behavior. Attention mechanisms are the key to this approach, providing a form of attention weights that can be used to explain why the model made a particular decision Xiong et al. (2022). Another type of intrinsic approach is to train the model to simultaneously perform the prediction task and generate the explanation for its predictions Fernandes et al. (2023). This “additional task” can be in the form of a text explanation or model prototype which embeds the semantic meaning of the prediction. However, the intrinsic approach is more difficult to apply because the user needs additional knowledge and understanding of the model’s architecture and inner workings.
5.0.3 Ethics and Fairness
@@
Deep learning models are increasingly being deployed in making high-stake decision including recruitment Freire and de Castro (2021), criminal justice Dass et al. (2023) and credit scoring Gicić et al. (2023). There are several advantages of deep learning-based systems in which, unlike humans, machines are able to process vast amounts of data and applications quickly and consistently. However, deep learning-based systems have the risk of being prone to biases present in the data used for training which can lead to unfairness and injustice. Numerous efforts have been made to mitigate this issue which can be categorized into modelling bias detection and modelling bias mitigation. Detection of modelling bias refers to the process of identifying and quantifying biases that may present in predictive models. This approach involves the use of statistical analysis, fairness metrics, counterfactual testing and human review to detect bias in the models. For instance, visualization-based methods such as attribution maps are used to indicate which regions are significant to the predictions Schaaf et al. (2021). This in turn can be used to detect and quantify bias using metrics such as Relevance Mass Accuracy, Relevance Rank Accuracy Accuracy and or Area over the perturbation curve (AOPC). In Giloni et al. (2022), two modules are presented for estimating bias in predictive models. The first module utilizes an unsupervised deep neural network with a custom loss function to generate hidden representation of the input data called bias vectors, revealing the underlying bias of each feature. The second module combines these bias vectors into a single vector representing the bias estimation of each feature, achieved by aggregating them using the absolute averaging operation. Bias mitigation refers to the process of reducing the presence of bias in predictive models, which can be done in three stages. The first stage combats bias by modifying the training data, either relabeling the labels or perturbing the feature values Iosifidis et al. (2019); Kehrenberg et al. (2020). The second stage addresses bias during the training of the model by applying regularization terms to the loss function to penalize discrimination. In Jain et al. (2023), a loss function based on bias parity score (BPS) is introduced to measure the degree of similarity of a statistical measure such as accuracy across different subgroups. The BPS term is added to the loss function as a regularizer to the original prediction task. The last stage mitigates bias after the predictive models have been successfully trained. This stage applies post-processing approaches such as reinforcement learning to obtain a fairer model Yang et al. (2023). For instance, the detection of minority classes is rewarded to prevent bias towards the majority class. This allows the model the generalize well across different patient demographics.
5.0.4 Lightweight Deep Learning Models
@@
Even though deep learning architectures have achieved state-of-the-art across various computer vision tasks, they often come with large model parameters Ige and Mohd Noor (2023). The architecture and complexity of a deep learning network determine the number of model parameters. The deeper the network, the larger the number of model parameters. However, deep learning models with large parameters often suffer limitations when deploying on end devices. For instance, a deep learning model developed for security monitoring by analyzing video data using 3D-CNNs might suffer deployment issues when deploying such models on low-resourced systems like smartphones or small-scale IoT devices. Model training and inference for deep learning models with large parameters demands substantial processing power. As the number of parameters increases, so does the computational complexity, resulting in longer model training duration and more hardware needs. Also, large parameter sizes translate to increased memory requirements, limiting their deployment on end devices. This is because these end devices often have battery, processor or memory capacity limitations. To address these challenges, it is important to develop sophisticated but lightweight architectures that can achieve state-of-the-art with few model parameters. Such lightweight models will be characterized by their ability to deliver competitive performance while mitigating computational complexity and memory requirements, making them well-suited for deployment on resource-constrained devices. An approach would be to develop novel lightweight plug-and-play modules that can be plugged to few layered deep learning architectures to improve feature learning without incurring additional model complexity. Other approaches could involve leveraging model compression techniques to reduce the size and computational complexity of deep learning models. Researchers can focus on improving pruning methods Li et al. (2019a), which can identify and eliminate redundant parameters or connections, thereby reducing the model’s footprint without compromising performance. Also, quantization techniques Yang et al. (2019) can be further explored to reduce the precision of weights and activations, therefore, enabling efficient representation with lower memory requirements. Also, knowledge distillation techniques Stanton et al. (2021) can be further investigated to facilitate the transfer of knowledge from a complex teacher model to a simpler student model, therefore, enabling compact yet effective representations. These areas are still open to contributions.
5.0.5 Adversarial Attack and Defense
@@
Adversarial attacks and defense mechanisms in deep learning represent a critical area of research and development, particularly as deep learning models become increasingly integrated into various applications. Adversarial attacks involves the deliberate manipulation of input data to mislead or deceive deep learning models, leading to incorrect predictions or behavior Akhtar and Mian (2018). Szegedy et al. Szegedy et al. (2013) was the first to identify this intriguing shortcoming of deep neural networks in image classification. They showed that even with their great accuracy, deep learning models are surprisingly vulnerable to adversarial attacks that take the form of tiny image changes that are (almost) invisible to human vision systems. A neural network classifier may radically alter its prediction about an image as a result of such an attack. Also, such a model can indicate high confidence in wrong predictions, which can be catastrophic for deep learning models deployed in medical or security fields, among many others. In generative models, several studies have investigated how adversarial attacks affect autoencoders and GANs, as seen in Tabacof et al. Tabacof et al. (2016) where a method to manipulate input images in a way that deceives variational autoencoders into reconstructing a totally different image was introduced. In recent times, the focus of adversarial attack research has been on images, but studies has shown that adversarial attacks are not limited to image data; they can also affect other types of data such as text, signals, audio, and video Zhang et al. (2020b); Jiang et al. (2019); Esmaeilpour et al. (2019). Future research can focus on exploring adversarial attacks in these domains and developing tailored defense mechanisms. Also, researchers can further investigate the practical implications of adversarial attacks in real-world scenarios, such as in autonomous vehicles, medical imaging, and cybersecurity. Understanding the potential impact of adversarial attacks in these applications can inform the development of more robust and secure systems.
6 Conclusions

### v5

@@ 5.1 Availability and Quality
Building and employing deep learning models face several challenges. The training of deep learning requires a large number of instances (examples) to achieve high accuracy and generalization [280]. Furthermore, the complexity of deep neural networks may lead to overfitting, where the model performs well on training data but fails to generalize on new, unseen data. This phenomenon frequently arises when the models is trained on insufficient data, highlighting the importance of diverse and extensive datasets. However, the data collection and annotation are time consuming and often require domain experts, specialized training and standardization [281]. Moreover, this process is prone to error and has the risk of introducing biases into the dataset which can significantly impact the performance of the trained model. One of the approaches to address this issue is transfer learning. Transfer learning involves the use of a deep learning model (known as pre-trained model) that is trained on a large dataset for solving a specific task (with a small dataset) [282]. The pre-trained model serves as a basis for the model training by fine-tuning the weights of the pre-trained model and adapting it to the new prediction task. This approach helps to mitigate the lack of training data in the target domain. Furthermore, transfer learning reduces computational resources required to train the model and helps faster convergence.
@@
Another approach that can be employed to address the lack of data is data augmentation. Data augmentation is a convenient method that increases the number of instances by performing transformation functions on the existing instances without changing the labels [283]. In the domain of computer vision, image transformation such as rotation, translation and cropping. However, it is important to consider the output of the transformation because the resultant may not represent the actual data. For example, flipping or adding noise to a signal may introduce distortion or changing the characteristics (trend, seasonality and cyclic variations) of the signal. Thus, careful consideration must be given to ensure that the generated instances still accurately represent the underlying patterns present in the data. Data augmentation can also be realized by generating synthetic data to supplement the training set. Synthetic data is artificially created data that resembles real data but is generated using statistical methods or deep generative models [284, 285]. The generated data can complement the less-diverse, limited datasets, providing a broader range of examples for the model to learn from. However, generating synthetic data that accurately reflects the characteristics of the real-world data is challenging. Careful consideration must be given to the choice of models and parameters used to ensure the synthetic data is realistic and representative of the real-world data.
5.2 Interpretability and Explainability
@@
Interpretability and explainability is crucial for building trust and understanding how predictive models make decisions especially in high-stake applications such as healthcare and medical image analysis [286]. However, as deep learning models become more intricate and complex with numerous layers, subnetworks and a large number of parameters, the models are often perceived as a “black box” and difficult to explain in terms of decision-making processes. Therefore, it is crucial for the researchers to focus on methods that provide insights into how a deep learning model performs the prediction and how its decisions are influenced by the input data, making it more transparent and trustworthy. Numerous methods have been proposed for interpreting and explaining the decisions of deep learning models which can be categorized into visualization (feature attribution), model distillation and intrinsic (explainable by itself). Visualization methods involve the use of scientific visualization such as saliency maps or heatmaps to express the explanation by highlighting the degree of association between the inputs and the predictions [287]. The heatmaps identify the saliency of the input features influencing the model’s predictions. The visualization approach is simple and intuitive and can be applied to tabular data and image data. Furthermore, it can be used to identify and debug issues in deep learning models, leading to improved performance and robustness.
@@
Model distillation is an approach to approximating a complex model by fitting a simpler model using the training set. The simpler model is built typically using a simpler or interpretable algorithm such as linear regression, decision tree or rule-based methods [288]. In this approach, the simpler model is trained to resemble the predictive behavior of the complex model. Then, the simpler model may serve as the proxy or surrogate model for explaining the complex model. Model distillation can be used together with visualization to further enhance the interpretability of the complex model [289]. Model distillation seeks explanations of the models that were never designed to be explainable. Ideally, the explanation of a deep learning model’s prediction should be included as part of the model output, or the explanation can be derived from the architecture of the model. This is because an intrinsic model can learn not only the mapping between the input and output, but also generate an explanation of the prediction that is faithful to the model’s behavior. Attention mechanisms are the key to this approach, providing a form of attention weights that can be used to explain why the model made a particular decision [290]. Another type of intrinsic approach is to train the model to simultaneously perform the prediction task and generate the explanation for its predictions [291]. This “additional task” can be in the form of a text explanation or model prototype which embeds the semantic meaning of the prediction. However, the intrinsic approach is more difficult to apply because the user needs additional knowledge and understanding of the model’s architecture and inner workings.
5.3 Ethics and Fairness
@@
Deep learning models are increasingly being deployed in making high-stake decision including recruitment [292], criminal justice [293] and credit scoring [294]. There are several advantages of deep learning-based systems in which, unlike humans, machines are able to process vast amounts of data and applications quickly and consistently. However, deep learning-based systems have the risk of being prone to biases present in the data used for training which can lead to unfairness and injustice. Numerous efforts have been made to mitigate this issue which can be categorized into modelling bias detection and modelling bias mitigation. Detection of modelling bias refers to the process of identifying and quantifying biases that may present in predictive models. This approach involves the use of statistical analysis, fairness metrics, counterfactual testing and human review to detect bias in the models. For instance, visualization-based methods such as attribution maps are used to indicate which regions are significant to the predictions [295]. This in turn can be used to detect and quantify bias using metrics such as Relevance Mass Accuracy, Relevance Rank Accuracy Accuracy and or Area over the perturbation curve (AOPC). In [296], two modules are presented for estimating bias in predictive models. The first module utilizes an unsupervised deep neural network with a custom loss function to generate hidden representation of the input data called bias vectors, revealing the underlying bias of each feature. The second module combines these bias vectors into a single vector representing the bias estimation of each feature, achieved by aggregating them using the absolute averaging operation. Bias mitigation refers to the process of reducing the presence of bias in predictive models, which can be done in three stages. The first stage combats bias by modifying the training data, either relabeling the labels or perturbing the feature values [297, 298]. The second stage addresses bias during the training of the model by applying regularization terms to the loss function to penalize discrimination. In [299], a loss function based on bias parity score (BPS) is introduced to measure the degree of similarity of a statistical measure such as accuracy across different subgroups. The BPS term is added to the loss function as a regularizer to the original prediction task. The last stage mitigates bias after the predictive models have been successfully trained. This stage applies post-processing approaches such as reinforcement learning to obtain a fairer model [300]. For instance, the detection of minority classes is rewarded to prevent bias towards the majority class. This allows the model the generalize well across different patient demographics.
5.4 Lightweight Deep Learning Models
@@
Even though deep learning architectures have achieved state-of-the-art across various computer vision tasks, they often come with large model parameters [197]. The architecture and complexity of a deep learning network determine the number of model parameters. The deeper the network, the larger the number of model parameters. However, deep learning models with large parameters often suffer limitations when deploying on end devices. For instance, a deep learning model developed for security monitoring by analyzing video data using 3D-CNNs might suffer deployment issues when deploying such models on low-resourced systems like smartphones or small-scale IoT devices. Model training and inference for deep learning models with large parameters demands substantial processing power. As the number of parameters increases, so does the computational complexity, resulting in longer model training duration and more hardware needs. Also, large parameter sizes translate to increased memory requirements, limiting their deployment on end devices. This is because these end devices often have battery, processor or memory capacity limitations. To address these challenges, it is important to develop sophisticated but lightweight architectures that can achieve state-of-the-art with few model parameters. Such lightweight models will be characterized by their ability to deliver competitive performance while mitigating computational complexity and memory requirements, making them well-suited for deployment on resource-constrained devices. An approach would be to develop novel lightweight plug-and-play modules that can be plugged to few layered deep learning architectures to improve feature learning without incurring additional model complexity. Other approaches could involve leveraging model compression techniques to reduce the size and computational complexity of deep learning models. Researchers can focus on improving pruning methods [301], which can identify and eliminate redundant parameters or connections, thereby reducing the model’s footprint without compromising performance. Also, quantization techniques [302] can be further explored to reduce the precision of weights and activations, therefore, enabling efficient representation with lower memory requirements. Also, knowledge distillation techniques [303] can be further investigated to facilitate the transfer of knowledge from a complex teacher model to a simpler student model, therefore, enabling compact yet effective representations. These areas are still open to contributions.
5.5 Adversarial Attack and Defense
@@
Adversarial attacks and defense mechanisms in deep learning represent a critical area of research and development, particularly as deep learning models become increasingly integrated into various applications. Adversarial attacks involves the deliberate manipulation of input data to mislead or deceive deep learning models, leading to incorrect predictions or behavior [304]. Szegedy et al. [305] was the first to identify this intriguing shortcoming of deep neural networks in image classification. They showed that even with their great accuracy, deep learning models are surprisingly vulnerable to adversarial attacks that take the form of tiny image changes that are (almost) invisible to human vision systems. A neural network classifier may radically alter its prediction about an image as a result of such an attack. Also, such a model can indicate high confidence in wrong predictions, which can be catastrophic for deep learning models deployed in medical or security fields, among many others. In generative models, several studies have investigated how adversarial attacks affect autoencoders and GANs, as seen in Tabacof et al. [306] where a method to manipulate input images in a way that deceives variational autoencoders into reconstructing a totally different image was introduced. In recent times, the focus of adversarial attack research has been on images, but studies has shown that adversarial attacks are not limited to image data; they can also affect other types of data such as text, signals, audio, and video [307, 308, 309].
6 Summary and Future Directions
We have discussed the state-of-the-art applications and challenges of deep learning in computer vision, natural language processing and time series analysis. In this section, we summarize the advancements made which can be grouped into model architecture, contextual enhancement, and loss function and optimization. Finally, we discuss possible future works in these domains.
Model Architecture: The recent advancements in computer vision focus on vision transformers (ViT) for scalable representation learning [18]. In image classification and image segmentation, ViTs are utilized to capture global context and relationships between different parts of an image, while in object detection, ViTs are used to streamline the detection process by eliminating the traditional components like anchor boxes and non-maximum suppression [carion_end--end_2020]. However, ViTs lack the ability to exploit local spatial features and struggle with hyperparameter sensitivity and performance on smaller datasets. To address this issue, researchers have developed hybrid deep learning architectures such as conformer [90] and MaxViT [91] that combine both transformers and convolutional neural networks to capture both local and global features. Similar trend is observed in modern natural language processing whereby the basis of the deep learning models is transformer architectures such as BERT [141] and its variants to achieve remarkable accuracy in tasks such as text classification, machine translation and text generation.
While recent advancements in computer vision and natural language processing have predominantly relied on transformer architectures, the field of time series analysis focuses on a different trajectory. Notably, areas such as human activity recognition, speech recognition and financial prediction have seen significant progress through the development of hybrid architectures that do not depend on transformers. Instead, these advancements leverage recurrent neural networks and convolutional neural networks, often incorporating attention mechanisms. The hybrid architectures allow the models to capture both local features and temporal dependencies in the data, which is crucial for time series analysis [196, ige2023deep]. Similar trends can be observed in ECG and EEG classification whereby not only convolutional and recurrent neural networks are used, but also generative models such as variational autoencoder [268] and generative adversarial network to generate data to address challenges posed by limited datasets and overfitting [252]. The generator of the generative model is composed of layers of BiLSTM, while the discriminator is composed of convolutional neural networks.
Contextual Enhancement: Recent advancements in text classification focus on capturing contextual information utilizing approaches such as joint embeddings of labels [148] and words as well as aspect-aware methods that enhance feature extraction [150]. Conversely, in machine translation, although transformers have proven effective, they often struggle to capture nuanced language features and contexts. To address this issue, various strategies have been introduced such as concatenating contextual sentences and context-sensitive training [158, 159]. In image generation, the researchers focus on improving the quality and relevance of generated images by enhancing the alignment between input text and images, while also addressing issues like fuzzy shapes and diversity through innovative deep learning architectures such as incorporating attention mechanisms and fusion modules into the model. Furthermore, recent models like DF-GAN [130] and DMF-GAN [131] showcase a shift towards single-stage generators with regularization strategies that maintain details while enhancing diversity. In text generation, BERT and its variants are the basis of the deep learning models. The research focuses on hierarchical architectures, capturing long text dependencies [171] and leveraging domain knowledge through label sequence [172] and conversation sequence [168] to bridge semantic gaps. Overall, the advancements show a clear trend toward improving contextual awareness and feature representation.
Loss Function and Optimization: Researchers are increasingly focused on refining loss functions to improve classification accuracy. For instance, in image classification, some studies introduce additive terms to the cross-entropy loss to reward well-classified instances [82], while others propose asymmetric polynomial loss functions that prioritize positive instances to tackle class imbalance [83] and adaptive loss function that dynamically adjust the weights assigned to class-level components based on model performance [310]. Another approach to address class imbalanced is data augmentation through generating virtual samples [311], oversampling with target-aware autoencoders for estimating target values for new features [312]. In time series analysis, similar approach has been proposed such as penalizing misclassification of minority classes [313], maximizing the minimum recall of the classes [314]. Improved training strategies have been proposed to address class imbalanced data such as iteratively selecting the most informative instances [315] and contrastive learning to keep the instances of each fine-grained clusters away from the minority class [316].
Future Directions: Looking ahead, the future of deep learning presents numerous opportunities for growth and innovation. Future models could explore non-sequential hybrid architectures of transformer and convolutional neural networks to leverage the strengths of both approaches, enhancing performance in image classification, object detection and image segmentation. Furthermore, researchers could investigate simpler or lightweight architectures and new training schemes to address the long training time associated with transformers. Future research in image generation may explore advanced frameworks or architectures that further integrate semantic understanding, perhaps by employing hierarchical attention modules and/or fusion modules to capture both local and global features more effectively. Additionally, incorporating unsupervised learning or self-supervised learning approaches could reduce reliance on labeled datasets, allowing models to learn from a diverse range of inputs. In the area of natural language processing, future research in text classification and machine translation could explore the integration of external knowledge bases and domain-specific embeddings may further improve context understanding and label alignment. Additionally, refining context incorporation methods such as dynamic context updating during translation may enhance the model performance.
Future research in image generation may explore advanced frameworks or architectures that further integrate semantic understanding, perhaps by employing hierarchical attention modules and/or fusion modules to capture both local and global features more effectively. Additionally, incorporating unsupervised learning or self-supervised learning approaches could reduce reliance on labeled datasets, allowing models to learn from a diverse range of inputs. In text generation, advancing towards more interactive and adaptive systems that can maintain context over extended dialogues and narratives is crucial. Furthermore, frameworks that can leverage multi-modalities (text, audio and visual) for richer contextual understanding may facilitate the development of more advanced applications across these domains. Future research can also focus on exploring adversarial attacks in these domains and developing tailored defence mechanisms. Also, researchers can further investigate the practical implications of adversarial attacks in real-world scenarios, such as in autonomous vehicles and medical imaging. Understanding the potential impact of adversarial attacks in these applications can inform the development of more robust and secure systems.
Due to the sensitive nature of financial research, future work can focus on enhancing the interpretability of deep learning models in financial predictions. Researchers should explore techniques to explain the predictions of models, to improve trust and understanding of model decisions, which is essential for adoption in human activity recognition, speech recognition and finance. Furthermore, future models could explore the use of transformers to enhance feature representation for time series analysis. Future research can also incorporate attention mechanisms into hybrid generative model architectures to improve the quality of generated signals. Additionally, real-time processing capabilities are paramount, future work can develop efficient algorithms for real-time processing of ECG and EEG signals. This could involve optimizing existing architectures and leveraging hardware acceleration techniques to enable real-time inference on resource-constrained devices such as wearable sensors and implantable devices. Future models could explore methods for adapting or personalizing models to account for inter-subject variability and improve performance on individual subjects. Also, EEG electrodes cover only a fraction of the brain’s surface, resulting in limited coverage of neural activity. Deep learning models could investigate strategies to infer activity from unobserved brain regions or integrate information from multiple modalities to provide more comprehensive coverage. These areas can still be further explored.
7 Conclusion